Please answer question 3
Find all (infinitely many) solutions of the system
Answer #1

3)

A.

given encryption function (ax+b)mod m = (13x+17)

mod m

Hence a=13 b=17 m=26

Decryption function = a^-1 *(x-b)mod m

Modular multiplicative inverse of a mod m is

i.e x = 13^-1(mod 26)

13x = 1(mod 26)

there is no inverse for above function since for any x the function doesnt give remainder one.

Hence the decryption is not possible.

B.

Given word = happy = 7 0 15 15 24

Apply y = (13x+17)mod 26

Then y0=91+17 mod 26=108 mod 26= 4

y1=17 mod 26=17

y2= 195+17 mod 26= 212 mod 26=4

y3= 195+17 mod 26=4

y4= 312+17 mod 26 =17

Hence y = 4 17 4 4 17